# Proper Learnability and the Role of Unlabeled Data

### v1

@@ 4 Obstructions to Characterizing Proper Learnability
We now direct our attention from the distribution-fixed PAC to the classical PAC model, and ask: Under what conditions is a learning problem properly learnable? Perhaps the most natural approach is to search for a combinatorial characterization of proper learnability, by analogy with existing characterizations of (improper) PAC learnability. We demonstrate several obstructions to any such approach. Together, they imply that proper learnability cannot be characterized by any property of finite character, as described in Ben-David et al. (2019). Our results also imply, more generally, that proper learnability cannot be characterized by any condition which is monotone, or which considers only “finite projections” of the hypothesis class. Throughout the section, we remain in the setting of multiclass classification in the (realizable) PAC model. Along with binary classification, this forms perhaps the most fundamental setting of supervised learning.
We now demonstrate a central technical result of the section: the pathological learning problem of EMX learning — which was not originally phrased as a supervised learning problem (Ben-David et al., 2019) — can in fact be witnessed as an instance of proper multiclass learning. We first recall the standard definition of EMX learning, and the result for which it was designed: EMX learnability can be logically undecidable.
@@
The learning problem associated to $\mathcal{F}$ is said to be EMX-learnable if there exists a learner which outputs sets of measure arbitrarily close to 1 as $|S|\to\infty$, with high probability over $S$ (and uniformly in the adversary’s choice of $P$). We will often abbreviate this by stating that $\mathcal{F}$ itself is, or is not, EMX-learnable. A breakthrough result of Ben-David et al. (2019) demonstrated that EMX learning can be undecidable, depending upon the cardinality of the underlying set $\mathcal{X}$.
Theorem 4.2 (Ben-David et al. (2019)).
@@
We now demonstrate that EMX learning can be witnessed within multiclass classification. Our proof employs techniques developed by Daniely et al. (2015) and Daniely and Shalev-Shwartz (2014) in their design of the first Cantor class.
@@
Then set =$\mathcal{H}=\{h_{A}:A\subseteq\mathcal{X},|\mathcal{X}\setminus A|<\infty\}$. That is, each $h_{A}\in\mathcal{H}$ outputs the label $\star$ on a finite set of points, and the label $A$ elsewhere. Note that the label $A$, in only being output by the hypothesis $h_{A}$, completely reveals the identity of $h_{A}$ as the true labeling function. In fact, if a learner ever observes any label other than $\star$ in the training set, then it has fully identified the true labeling function. Then in order to learn $\mathcal{H}$, one need only consider learnability with respect to pairs $(D,h)$ where $D$ places full measure on $h^{-1}(\star)$. In particular, upon observing a training set $S$ with $|S|=n$, either $S$ contains a non-$\star$ label, rendering learning trivial, or $S$ only contains the $\star$ label, from which one can conclude that there is $o_{n}(1)$ probability of ever observing a non-$\star$ label (and thus learning reduces to correctly predicting the $\star$ labels). [^3] More precisely, for any distribution $D$ and sample size $n$, either $D$ places $O(\frac{1}{n})$ mass on non-$\star$ labels (thus learning reduces to correctly predicting the $\star$ label), or $S\sim D^{n}$ contains a non-$\star$ label with probability $1-o(1)$.
[…]
@@
An immediate consequence of Theorem 4.5 is that proper multiclass learnability is not a property of finite character, in the sense of Ben-David et al. (2019). This stands in stark contrast to the existing characterization of improper learnability by the DS dimension, and to similar dimension-based results across supervised learning. We now demonstrate that proper learnability is furthermore not a local property. That is, there exist hypothesis classes sharing all local behaviors yet differing in their proper learnability.
@@
Let $\mathcal{F}$ be a set of cardinality $|\mathcal{F}|\geq\aleph_{\omega}$. Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be the multiclass problem associated to $\mathcal{F}$, as per Proposition 4.3. By Theorem 4.2, $\mathcal{F}$ is not EMX-learnable and thus $\mathcal{H}$ is not properly learnable. Let $f^{\star}\colon\mathcal{X}\to\mathcal{Y}$ be the constant function which only outputs the $\star$ label. Then =$\mathcal{H}^{\prime}=\mathcal{H}\cup\{f^{\star}\}$ is certainly properly learnable, by the learner which outputs $f^{\star}$ when it only sees the $\star$ label in the training set, and otherwise sees a label $A\neq\star\in\mathcal{Y}$ which fully identifies the true labeling function as $h_{A}$. Yet for any finite $S\subseteq\mathcal{X}$, we have that $\mathcal{H}|_{S}=\mathcal{H}^{\prime}|_{S}$, because the behavior $f^{*}|_{S}$ appears in $\mathcal{H}|_{S}$ as the restriction of (for instance) $h_{\mathcal{X}\setminus S}$ to $S$. This completes the argument. ∎
@@ Appendix A Omitted proofs
Inequality (1) follows from the standard leave-one-out argument of transductive learning, which establishes that any learner $A$ incurring transductive error $\leq\epsilon$ on samples of size $n$ automatically incurs expected error $\leq\epsilon$ as well (Haussler et al., 1994). More explicitly, for any such learner $A$ and realizable distribution $D$ over $\mathcal{X}\times\mathcal{Y}$, one has:
@@
The third line makes use of the fact that a sample $T\sim\operatorname{Unif}(S)^{m-1}$ avoids any given point $x_{i}\in S$ with probability $\geq\frac{1}{e}$. (Recall that a transductive learning instead on a labeled dataset $|S|=m$ employs training sets of size $m-1$.) In particular, let $f(m)=(1-\frac{1}{m})^{m-1}$ denote the probability of a given $x_{i}\in S$ being avoided by $T\sim\operatorname{Unif}(S)^{m-1}$. First note that
@@
as $\left(\frac{x-1}{x}\right)^{x}>0$, $x-1>0$, and

### v2

@@ 4 Obstructions to Characterizing Proper Learnability
We now direct our attention from the distribution-fixed PAC to the classical PAC model, and ask: Under what conditions is a learning problem properly learnable? Perhaps the most natural approach is to search for a combinatorial characterization of proper learnability, by analogy with existing characterizations of (improper) PAC learnability. We demonstrate several obstructions to any such approach. Together, they imply that proper learnability cannot be characterized by any property of finite character, as described in ben2019learnability. Our results also imply, more generally, that proper learnability cannot be characterized by any condition which is monotone, or which considers only “finite projections” of the hypothesis class. Throughout the section, we remain in the setting of multiclass classification in the (realizable) PAC model. Along with binary classification, this forms perhaps the most fundamental setting of supervised learning.
We now demonstrate a central technical result of the section: the pathological learning problem of EMX learning — which was not originally phrased as a supervised learning problem (ben2019learnability) — can in fact be witnessed as an instance of proper multiclass learning. We first recall the standard definition of EMX learning, and the result for which it was designed: EMX learnability can be logically undecidable.
@@
The learning problem associated to $\mathcal{F}$ is said to be EMX-learnable if there exists a learner which outputs sets of measure arbitrarily close to 1 as $|S|\to\infty$, with high probability over $S$ (and uniformly in the adversary’s choice of $P$). We will often abbreviate this by stating that $\mathcal{F}$ itself is, or is not, EMX-learnable. A breakthrough result of ben2019learnability demonstrated that EMX learning can be undecidable, depending upon the cardinality of the underlying set $\mathcal{X}$.
Theorem 4.2 (ben2019learnability).
@@
We now demonstrate that EMX learning can be witnessed within multiclass classification. Our proof employs techniques developed by daniely2015multiclass and DS14 in their design of the first Cantor class.
@@
Then set $\mathcal{H}=\{h_{A}:A\subseteq\mathcal{X},|\mathcal{X}\setminus A|<\infty\}$. That is, each $h_{A}\in\mathcal{H}$ outputs the label $\star$ on a finite set of points, and the label $A$ elsewhere. Note that the label $A$, in only being output by the hypothesis $h_{A}$, completely reveals the identity of $h_{A}$ as the true labeling function. In fact, if a learner ever observes any label other than $\star$ in the training set, then it has fully identified the true labeling function. Then in order to learn $\mathcal{H}$, one need only consider learnability with respect to pairs $(D,h)$ where $D$ places full measure on $h^{-1}(\star)$. In particular, upon observing a training set $S$ with $|S|=n$, either $S$ contains a non-$\star$ label, rendering learning trivial, or $S$ only contains the $\star$ label, from which one can conclude that there is $o_{n}(1)$ probability of ever observing a non-$\star$ label (and thus learning reduces to correctly predicting the $\star$ labels). [^3] More precisely, for any distribution $D$ and sample size $n$, either $D$ places $O(\frac{1}{n})$ mass on non-$\star$ labels (thus learning reduces to correctly predicting the $\star$ label), or $S\sim D^{n}$ contains a non-$\star$ label with probability 1o(1)1-o(1).
Now suppose that $\mathcal{H}$ is properly learnable. Then there exists a learner $\mathcal{A}$ for $\mathcal{H}$ with the following property: for any marginal distribution $D$ over $\mathcal{X}$ with finite support, when $\mathcal{A}$ observes a training sample $S=(x_{i},\star)_{i\in[n]}$ with $x_{i}\overset{\text{i.i.d.\ }}{\sim}D$, it emits a hypothesis $h_{A}\in\mathcal{H}$ such that $\mathcal{X}\setminus A$ is finite and has large $D$-measure. By modifying $\mathcal{A}$ to receive only the data of $(x_{i})_{i\in[n]}$ and to emit $\mathcal{X}\setminus A$ rather than $h_{A}$, we produce an EMX learner for $\mathcal{F}$. Conversely, any EMX learner $\mathcal{A}$ for $\mathcal{F}$ gives rise to a PAC learner for $\mathcal{H}$ by nearly identical reasoning, i.e., by reformatting its input and output. ∎
@@
An immediate consequence of Theorem 4.5 is that proper multiclass learnability is not a property of finite character, in the sense of ben2019learnability. This stands in stark contrast to the existing characterization of improper learnability by the DS dimension, and to similar dimension-based results across supervised learning. We now demonstrate that proper learnability is furthermore not a local property. That is, there exist hypothesis classes sharing all local behaviors yet differing in their proper learnability.
@@
Let $\mathcal{F}$ be a set of cardinality $|\mathcal{F}|\geq\aleph_{\omega}$. Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be the multiclass problem associated to $\mathcal{F}$, as per Proposition 4.3. By Theorem 4.2, $\mathcal{F}$ is not EMX-learnable and thus $\mathcal{H}$ is not properly learnable. Let $f^{\star}\colon\mathcal{X}\to\mathcal{Y}$ be the constant function which only outputs the $\star$ label. Then $\mathcal{H}^{\prime}=\mathcal{H}\cup\{f^{\star}\}$ is certainly properly learnable, by the learner which outputs $f^{\star}$ when it only sees the $\star$ label in the training set, and otherwise sees a label $A\neq\star\in\mathcal{Y}$ which fully identifies the true labeling function as $h_{A}$. Yet for any finite $S\subseteq\mathcal{X}$, we have that $\mathcal{H}|_{S}=\mathcal{H}^{\prime}|_{S}$, because the behavior $f^{*}|_{S}$ appears in $\mathcal{H}|_{S}$ as the restriction of (for instance) $h_{\mathcal{X}\setminus S}$ to $S$. This completes the argument. ∎
@@ Appendix A Omitted proofs
Inequality (1) follows from the standard leave-one-out argument of transductive learning, which establishes that any learner $A$ incurring transductive error $\leq\epsilon$ on samples of size $n$ automatically incurs expected error $\leq\epsilon$ as well (haussler1994predicting). More explicitly, for any such learner $A$ and realizable distribution $D$ over $\mathcal{X}\times\mathcal{Y}$, one has:
@@
The third line makes use of the fact that a sample $T\sim\operatorname{Unif}(S)^{m-1}$ avoids any given point $x_{i}\in S$ with probability $\geq\frac{1}{e}$. (Recall that a transductive learning instead on a labeled dataset $|S|=m$ employs training sets of size m1m-1.) In particular, let $f(m)=(1-\frac{1}{m})^{m-1}$ denote the probability of a given $x_{i}\in S$ being avoided by $T\sim\operatorname{Unif}(S)^{m-1}$. First note that
@@
as $\left(\frac{x-1}{x}\right)^{x}>0$, x1>0x-1>0, and